OpenStudy (anonymous):

3\frac{2}{9}\; \div \; 1\frac{2}{27} Describe the process for solving the problem step by step. Remember to show your work for each step. Don't forget to simplify the answer

OpenStudy (gabebae):

Again.. for the second time I can't understand this

OpenStudy (anonymous):

\[3\frac{2}{9}\; \div \; 1\frac{2}{27} \]

OpenStudy (gabebae):

Alright

OpenStudy (anonymous):

U see it?

OpenStudy (gabebae):

Turn both fractions into an improper fraction

OpenStudy (anonymous):

How? XD

OpenStudy (anonymous):

3x2+9?

OpenStudy (gabebae):

\[\frac{ 11 }{ 5 }\] is an improper fraction

OpenStudy (gabebae):

no it's 9x3 + 2

OpenStudy (gabebae):

You were close..

OpenStudy (anonymous):

27+2= 29. 29/9

OpenStudy (anonymous):

Very Close...

OpenStudy (anonymous):

Is 29/9 the first frac?

OpenStudy (gabebae):

Hold on

OpenStudy (anonymous):

29/9 / 29/9?

OpenStudy (gabebae):

\[\frac{ 29 }{ 9 } \div \frac{ 29 }{ 27 }\]

OpenStudy (gabebae):

Those are the improper fractions

OpenStudy (anonymous):

1/243???? @GabeBae

OpenStudy (gabebae):

Where/how did you get that from?

OpenStudy (anonymous):

I don't know

OpenStudy (gabebae):

It's actually much easier than you think :')

OpenStudy (anonymous):

XD. I have no idea what the answer is

OpenStudy (gabebae):

\[\frac{ 29 }{ 9 } \div \frac{ 29 }{ 27 }\] You would flip the second fraction and find the reciprocal and that would turn it into a multiplication sentence This only works when you're dividing So then you would get \[\frac{ 29 }{ 9 } \times \frac{ 27 }{ 29 }\]

OpenStudy (gabebae):

Then you would just multiply the numerator and denominator

OpenStudy (anonymous):

3?

OpenStudy (gabebae):

29x29 for the numerator (top number of the fraction) 9x29 for the denominator (bottom number of the fraction)

OpenStudy (gabebae):

\[\frac{ 841 }{ 261 }\] Would be your answer but you need to simplify it

OpenStudy (anonymous):

Would the answer be 3 than?

OpenStudy (gabebae):

No..

OpenStudy (anonymous):

ok..hmm

OpenStudy (anonymous):

3 20/87?

OpenStudy (anonymous):

Nvm, that wouldn't be it

OpenStudy (anonymous):

I don't knwo what to simplyify by? What about 2 or 4?

OpenStudy (gabebae):

I was eating dinner

OpenStudy (anonymous):

o

OpenStudy (anonymous):

Yeah I'm stupid with fractions and everything is like .... I hate my life XD

OpenStudy (gabebae):

Yup that would be your answer as a fraction

OpenStudy (gabebae):

I need my privacy with food When it comes to food it's like locking yourself in your room and making out with it

OpenStudy (gabebae):

I ate my love

OpenStudy (anonymous):

lol

OpenStudy (gabebae):

Do you need more help or are you good now?

OpenStudy (anonymous):

I have 2 open ended and science...then I am done. yeah sue me pls. Then that might get me out of school work that;s due by midnight tonight :D

OpenStudy (gabebae):

Are you homeschooled?

OpenStudy (anonymous):

Yeppers

OpenStudy (gabebae):

Same..

OpenStudy (anonymous):

XD

OpenStudy (gabebae):

Do we get a day off tomorrow?

OpenStudy (anonymous):

I don't know. I didn't have class connects and the teachers didn't say anything.

OpenStudy (gabebae):

Show me the two open ended questions

OpenStudy (gabebae):

Let's get this pellet over with

OpenStudy (gabebae):

Again.. it was blurred out :')

OpenStudy (anonymous):

How would I explain ti step by step though/

OpenStudy (anonymous):

Should*

OpenStudy (gabebae):

Wait was this the open ended question?

OpenStudy (anonymous):

Yes.....2 of them and the last one. I nee dto explain it XD...

OpenStudy (gabebae):

I know this might be random but listen to this song https://www.youtube.com/watch?v=RGk68TkIeFI

OpenStudy (gabebae):

The eargasm.. is too much </3 SENPAI

OpenStudy (anonymous):

That's really good music 0-0

OpenStudy (gabebae):

Yup this is the type of music I listen to

OpenStudy (gabebae):

I get these songs from monstercat

OpenStudy (gabebae):

or NoCopyRightSounds

OpenStudy (anonymous):

lol

OpenStudy (gabebae):

Alright show me the last question I have to go soon

OpenStudy (anonymous):

o. can u stay for like 30 more min? XD 2 more open ended + science X:

OpenStudy (anonymous):

Yep, I'm all out of luck today XD

OpenStudy (gabebae):

Meh show me the questions?

OpenStudy (gabebae):

How long is the science?

OpenStudy (anonymous):

32 questions. but most are mutliple choice and should be quick. 4 open ended. but I know 2 of them

OpenStudy (anonymous):

And I know some of the science like half of it too

OpenStudy (gabebae):

What type of science?

OpenStudy (anonymous):

Astronomy

OpenStudy (gabebae):

Damn.. I love that S H I T

OpenStudy (gabebae):

Science first? </3

OpenStudy (anonymous):

XD fine.

OpenStudy (gabebae):

:D

OpenStudy (anonymous):

Scientists use the term _________________ to describe how the force of gravity pulls on an object. A. Weight B. Mass C. Acceleration D. Friction

OpenStudy (anonymous):

A or B

OpenStudy (gabebae):

A

OpenStudy (anonymous):

What is the term that describes the mutual attraction of objects that have mass? A. Density B. The Big Bang C. Centrifugal force D. Gravity

OpenStudy (gabebae):

D

OpenStudy (anonymous):

How does mass affect gravity? A. The less mass an object has, the stronger the gravitational pull B. The more mass an object has, the stronger the gravitational pull C. The more mass an object has, the weaker the gravitational pull D. There is no relationship

OpenStudy (gabebae):

I would go with D Mass doesn't matter

OpenStudy (gabebae):

Damn I have to go I'm so sorry

OpenStudy (anonymous):

o..

OpenStudy (gabebae):

continue?

OpenStudy (anonymous):

@GabeBae

OpenStudy (gabebae):

mhm?